OpenStudy (anonymous):

This one is a little bit tricky: The surface area of a closed cylinder is given by the formula SA = 2(pi)r(r+h) where 'r' centimetres is the radius of the can and 'h' cm is the height. The height of a can of wood finish is 7 cm and its surface area is 231cm^2. (A) Use the quadratic formula to solve the equation and hence find the radius of the can

Directrix (directrix):

SA = 2*pi *r * h +2* pi r^2 231 = 2 * pi * r * 7 + 2 * pi * r^2

Directrix (directrix):

231 = 2 * π * r * 7 + 2 *π * r² 2 *π * r² + 14 π * r - 231 = 0

Directrix (directrix):

r = -10.501 or r = 3.5011 Check this.

Directrix (directrix):

@Chad123

OpenStudy (anonymous):

Your a genius. Thanks for your help

OpenStudy (anonymous):

How did you get 3.5?

OpenStudy (anonymous):

Could you please show me how? Thx

Directrix (directrix):

@Chad123 I am not a genius. You can use the quadratic formula to crank out the values of r. Follow this link to see where I did it. http://www.wolframalpha.com/input/?i=+2+*%CF%80+*+r%C2%B2++%2B+14+%CF%80+*+r++-+231+%3D+0

Directrix (directrix):

Scroll down and click on "approximate form."